OpenStudy (anonymous):

help! solve.. x+2 3-x 2 ____ + _____ = ___ 15 5 3

OpenStudy (jdoe0001):

$$\large { \cfrac{x+2}{15}+\cfrac{3-x}{5}=\cfrac{2}{3} } $$ if you were to multiply both sides by "15", what would it give you?

OpenStudy (phi):

jdoe gave you a good idea. multiply both sides (all terms) by 15 like this: \[ 15 \cdot \cfrac{x+2}{15}+15 \cdot\cfrac{3-x}{5}=15 \cdot\cfrac{2}{3} \] can you simplify that ?

OpenStudy (anonymous):